OpenStudy (anonymous):

use mental math for 8 times 6

OpenStudy (amistre64):

what, pray tell, is mental math?

OpenStudy (anonymous):

lol I was wonderin the same thing

OpenStudy (anonymous):

how do you work on it with mental math i need help with that

OpenStudy (anonymous):

ummm add 8 six times

OpenStudy (anonymous):

well that would work

OpenStudy (shadowlegendx):

Mental math is the use of your brain to do mathematical calculations

OpenStudy (anonymous):

yeah, or add 8x3=24 twice

OpenStudy (shadowlegendx):

So for this question, you would have to do \[8 \times 6\] In your head

OpenStudy (amistre64):

we used to call that memorizing our times tables then

OpenStudy (lovelyharmonics):

use you multiplication tables. get flash cards. write it down 100x. whatever works best for you in order to memorize the material.

OpenStudy (anonymous):

yes Google times table and study it!

OpenStudy (anonymous):

thank you it is just my teacher expects more i know my multiplication it is just that i thought there is a another way

OpenStudy (amistre64):

8 times (5+1), mentally is 40+8

OpenStudy (lovelyharmonics):

my teacher use to print off worksheets of all the times tables on them and say "you have a minute to finish this, go!'" *everyone frantically flips the sheet over trying to complete it in time...

OpenStudy (anonymous):

you should know your simple multiplication facts

OpenStudy (anonymous):

i do

OpenStudy (lovelyharmonics):

that makes you learn it real fast when its a pass or fail grade.

OpenStudy (anonymous):

in your head

OpenStudy (anonymous):

my try is 10*6-(2*6)=60-6-6>.<

OpenStudy (anonymous):

@BSwan wayyy too complicated

OpenStudy (anonymous):

What is life... he's a bigger troll than me.

OpenStudy (lovelyharmonics):

oh my gosh.... there are 3 animal profile pics on this one question <3

OpenStudy (bibby):

yeah I mean subtracting 12 from 60, how will my puny brain handle that?

OpenStudy (anonymous):

eshhh >.< its not hard to remember :o

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

@lovelyharmonics I like bunnies

OpenStudy (anonymous):

8*6 itself already mentaly

OpenStudy (rational):

2^3*2*3 2^4*3 16*3 48

OpenStudy (squirrels):

My father would make me repeat my multiplication tables by answering each question out loud 10 times, and if I get one wrong I would get WHIPPED. Cried a lot but it works like a charm.

OpenStudy (bibby):

hehe reminds me of the good times when my brother and I "bonded" @squirrels

OpenStudy (anonymous):

@Squirrels my mom did something similar

OpenStudy (anonymous):

For this type of math, you get mental first and try to find \(8 \times 6\) in that very condition.. :P

OpenStudy (anonymous):

Bibs ♥

OpenStudy (kainui):

Pull a 2 out of the 6 and push it into the 8 to get 16*3 then separate it out: (10+6)*3 distribute 30+18 48 I don't know just making crap up tbh

OpenStudy (lovelyharmonics):

i had a friend who was repeatedly beaten using a wire hanger.... i guess that fits in... just minus the math

OpenStudy (bibby):

Is this the sexual abuse help line/group chat/support group

OpenStudy (anonymous):

This is entertaining

OpenStudy (bibby):

My name is bibby and I'm an alcoholic

OpenStudy (kainui):

My name is Kainui and I don't want to violate the code of conduct by saying what I am.

OpenStudy (dan815):

this is just mental

OpenStudy (anonymous):

\[6 \times 8 \implies 6 \times (10-2) \implies 60 - 12 \implies \color{green}{48}\]

OpenStudy (anonymous):

LOL Dan xD

OpenStudy (anonymous):

well we wont forget , just count the sticks :D 

OpenStudy (anonymous):

My name is stupid and I'm am stupid

OpenStudy (dan815):

lool kainui

OpenStudy (rational):

just for a variety : 8*6 (10-2)(10+2)/2 (10^2-2^2)/2 96/2 48

OpenStudy (lovelyharmonics):

oh my...

OpenStudy (rational):

good to see you back @lovelyharmonics :)

OpenStudy (anonymous):

waitt, 4x6\[4x \frac{ 4 }{ 8 }=1/2 4x, 2x, x=idk\]